A square loop of wire, of length A = 0.1 m on each side, has a mass of 50 g and pivot


1
Expert's answer
2021-11-04T10:16:34-0400

For the magnetic field B directed at angle θ to the plane of the loop, the torque around the axis is


"\tau=Il^2B\cos\theta."